OpenStudy (anonymous):

How do you change a number into a radical? Example: Square root of 24.

OpenStudy (rogue):

You factor 24 into 2 #s that when multiplied, give you 24. One of them has to be a perfect square. For 24, the 2 #s are 4 and 6.\[\sqrt {24} = \sqrt {4 *6} = \sqrt {4} * \sqrt {6} = 2 \sqrt {6}\]

OpenStudy (anonymous):

Thank you so much!

OpenStudy (rogue):

No problem, just try a few problems on your own and you should get the hang of it :)